OpenStudy (anonymous):

The lengths of the leg of a right triangle opposite an angle with measure of 30 degrees is 10. Find the length of its hypotenuse. a) 5 b) 10 c) 5 square root of 3 d) 20

OpenStudy (anonymous):

20

Parth (parthkohli):

The angle opposite to 30 is x and the hypotenuse is 2x. So, if x = 10, then hypotenuse = 2x = 20

OpenStudy (anonymous):

tan 30 =10/(the other side) so the other side =10/tan 30=10sqrt3 so hypotenuse =sqrt( 10^2 +10^2*3)=sqrt 400=20

OpenStudy (anonymous):

@ParthKohli has an easier solution